Oceania Limited acquired 100% of the share capital of Broadwater Limited for $600 000 in cash. Broadwater had total shareholder’s equity of $500 000. The book values of BroadwaterLimited’s assets were: buildings $300 000, machinery $180 000. The fair values of these assets were: buildings $360 000, machinery $200 000. The tax rate is 30%. The goodwill or bargain purchase0 is:

a. ​$20 000

b. ​$44 000

c. ​$180 000

d. ​$(44 000)
Answer #1

Other assets = total shareholder’s equity - Building - machinery

= 500000 - 300000 -180000 = $20,000

Goodwill = Purchase price - Fair value of land & building - other assets

= 600000 - 360000 - 200000 - 20000

= $20,000

Option a. is correct answer.